# Is Large-scale Pretraining the Secret to Good Domain Generalization?

### v2

@@ 5.1 Training and Evaluation Protocol
We aim to evaluate how well Domain Generalization (DG) methods perform across both pretraining-aligned(IP) and pretraining-misaligned(OOP) data. We adhere to the DomainBed evaluation methodology, where one domain is chosen as the target, and the remaining domains act as source domains. To maintain a sufficient amount of training data, we train all DG methods on the original, unsplit datasets. We use hyper-parameter values recommended by the original implementation authors for each method.
@@
After training, models are evaluated separately on the IP and OOP subsets, as well as the original, unsplit test domain. This approach allows us to measure how well each method generalizes to both pretraining-aligned and pretraining-misaligned data.
We follow the literature’s standard practice of computing performance per target domain and averaging the results across all domains. This method ensures that any domain imbalances do not disproportionately influence the final performance metrics.
Table 2: Benchmarking DG methods on DB-IP/OOP
@@
(a) Samples with high AlignmentScore values, indicating good alignment between image and text features learned during pretraining.
@@
5.2 Algorithms
Explicit Regularization towards Pretraining: Several recent DG methods leverage explicit regularization towards the initialization. These methods generally operate either in weight space (by regularizing or freezing model parameters) or in feature space (by aligning internal feature representations with those of the pretrained model).
  • MIRO (Cha et al., 2022): Minimizes the Mutual Information between DG model intermediate features and CLIP intermediate features.
  • Attention Tuning (Teterwak et al., 2023; Touvron et al., 2022): Freezes all parameters except those in the Multiheaded-Attention Layers.
  • VL2V-SD (Addepalli et al., 2024): Self-distills a linear combination of CLIP vision and text outputs into a model.
  • CLIPood (Shu et al., 2023): Regularizes both weights and output features. Weights are averaged between the pre-trained CLIP model and the fine-tuned model, and outputs are regularized using a loss that incorporates class information from the pre-trained text encoder.
  • Linear Probe - Fine Tuning (LP-FT) (Kumar et al., 2021): Freezes the backbone, trains a linear probe, and then performs full finetuning. An untrained linear probe can cause finetuning to update the frozen backbone needlessly, potentially unlearning discriminative features. This biases the model towards pre-trained weights with smaller gradient updates.
Flat Optima: Studies by Izmailov et al. (2018) and Cha et al. (2021) have shown that flat minima generalize better than sharp minima, as they make loss values less sensitive to perturbations in the loss surface, resulting in smaller increases in loss during domain shifts.
  • SWAD (Cha et al., 2021): Averages model parameters during training, determining the interval over which to average using validation loss over source domain.
  • Model Parameter Averaging (MPA) (Arpit et al., 2022): Starts averaging model parameters after a number of burn-in steps to find flat minima.
  • SAGM (Wang et al., 2023): An optimizer that explicitly optimizes for flat minima.
Baseline and Oracle: We also train a baseline and oracle model for lower-bound and upper-bound reference. The baseline model is an Empirical Risk Minimization (ERM) model that is finetuned on source domains and evaluated on target domains, and has been found to be effective for the DG task (Gulrajani & Lopez-Paz, 2020). The oracle model is trained on an 80% training split of all domains and evaluated on a 20% test split. The oracle model removes the OOD aspect of generalization and provides a reasonable upper bound for DG methods.
5.3 Results
We show results in Table 2(c) for DomainBed-OOP, DomainBed-IP, and standard DomainBed datasets. Underlined results represent the best performance of any single method (excluding method combinations), while the bold numbers show the highest performance overall, excluding the upper bound.
DG methods have nearly solved DomainBed-IP: In most datasets within DomainBed-IP, domain generalization approaches have nearly solved the problem of generalization. On three out of five datasets (DomainNet, OfficeHome, and PACS), the best DG method even outperforms the oracle! A notable exception is TerraIncognita, where CLIPood scores only 63.5%, far below the oracle’s 93.4%, highlighting this dataset as a challenging outlier. Interestingly, all three datasets where performance exceeds the oracle have an average IP AlignmentScore of 0.28, while the others have a lower average AlignmentScore of around 0.26. Therefore, the underperformance of TerraIncognita may be partially explained by its lower IP AlignmentScore, suggesting that alignment plays a significant role in DG performance, even within the IP case.
DG Methods leave much to be desired on DomainBed-OOP, but are still stronger than ERM: In DomainBed-OOP, we observe that even the top-performing DG methods struggle with low-alignment data. For example, CLIPood achieves 57.1% accuracy, which is a significant drop compared to its performance on DomainBed-IP (84.7%) and DomainBed-All (78.1%). Despite this, DG methods still outperform Empirical Risk Minimization (ERM), which scores only 47.0% on DomainBed-OOP. Therefore DG methods are better equipped to handle domain shifts than ERM, possibly due to weak transfer of knowlegdge from pre-trained features. Nevertheless, there is still substantial room for improvement on low-alignment samples.
@@ 5.3 Results
State-of-the-Art methods do not consistently outperform older methods on OOP data: While CLIPood (Shu et al., 2023) clearly outperforms other methods on DomainBed-All with 78.1% accuracy, it performs comparably to older methods on DomainBed-OOP. For instance, MIRO + MPA (Cha et al., 2022; Arpit et al., 2022) achieves 57.2% on DomainBed-OOP, which is nearly the same as CLIPood’s 57.1%. This suggests that CLIPood’s primary advantage comes from well-aligned samples in DomainBed-IP, where it reaches 84.7% accuracy, half a percent better than the next-best method.
@@
Model Parameter Averaging (MPA) boosts performance on OOP data: MPA performs significantly better than ERM on DomainBed-OOP, achieving 53.0% compared to ERM’s 47.0%. This is half a percent better than MIRO on OOP data, despite being 2% worse than MIRO on IP data. When combined with MIRO, MPA delivers the best performance on DomainBed-OOP, scoring 57.2%, slightly surpassing CLIPood. This suggests that MPA can complement other regularization-based methods like MIRO. On DomainBed-IP, MIRO + MPA scores 84.1%, close to CLIPood’s 84.7%, demonstrating versatility across both high- and low-alignment data. Interestingly, SWAD underperforms MPA on DomainBed-OOP by 2%, despite being conceptually similar. We attribute this to selecting the averaging interval on the source data, which introduces overfitting to source domains.
@@ 6 Conclusion
In this paper, we systematically explore how Domain Generalization (DG) methods rely on pre-trained feature alignment from models like CLIP. We hypothesize that the alignment between image and text embeddings during pre-training strongly predicts DG performance. Our experiments confirm this, showing that methods perform well on high-alignment samples (DomainBed-IP) but struggle on low-alignment data (DomainBed-OOP). While we focus on multi-source DG, we expect similar results in single-source finetuning due to the foundational role of pre-trained alignment. Notably, state-of-the-art methods like CLIPood perform near oracle-level on aligned data but see significant drops on misaligned samples. This suggests current DG methods rely on pre-trained features and fail to learn new, generalizable features from source domains. Moving forward, two paths emerge: developing DG methods that better learn generalizable features, or focusing on improving pre-trained backbones. While foundation models will continue to advance, there will always be specialized distributions where they fail. We hope these findings inspire further research into improving generalization on low-alignment data, pushing DG beyond reliance on pre-trained alignment.
@@ A.1 Training and Evaluation Details
Code is zipped in the supplementary.
We use a slightly modified MIRO Cha et al. (2022) codebase for training and evaluation. We use leave-one-out evaluation, where a model is trained on all domains except the evaluation domain. We emphasize that we use DomainBed-IP and DomainBed-OOP as evaluation data only, models are trained on full datasets.
For training, we use an OpenCLIP-ViT-B/16 Ilharco et al. (2021) trained on LAION-400M Schuhmann et al. (2021). We use default hyper-parameters as defined by  Cha et al. (2022). This includes a learning rate of 5e-5, weight decay of 0.0, a batch size of 32 per-domain, an Adam Optimizer, and no dropout for all methods.
@@
For benchmarked methods, we also use hyper-parameters found to be best in respective papers. For SWAD, we use an optimum patience parameter value of 3, overfit patience parameter value of 6, and tolerance ratio of 6. For MIRO, we use use regularizer loss weight of 1.0. For CORAL, we use a CORAL regularizer weight of 1.0, following  Cha et al. (2021). For LP-FT, we train the linear probe for 600 steps before unlocking the full backbone. For Model Parameter Averaging, we burn in the training for 600 steps before averaging iterates. For VL2V-SD and CLIPood, we directly use the author’s implementation and hyper-parameters, except initializing with OpenCLIP Ilharco et al. (2021).
@@ B.1 AlignmentScore vs Accuracy
In Figure 8, we plot all benchmarked methods from the main paper, with x-axis corresponding to AlignmentScore, and the y-axis corresponding to the Top-1 Accuracy. We normalize for dataset size, so that no dataset dominates the count. In Figures 9 through 13, we plot these statistics independently per dataset, and find the trends consistent across datasets.
@@ B.2 Per-Dataset Benchmarking Results
We expand Table 2(c) in the main paper into per-dataset results in Table 5 through 19.
@@ B.3 Similarity of Target to Pre-training
To evaluate the Image Similarity Hypothesis, we retrieve the nearest neighbors from the Laion-400M dataset Schuhmann et al. (2021). This raises the question of how similar the target domains are to the pre-training data and whether the source domains might be even more similar. To investigate this, we compute Maximum Mean Discrepancy (MMD) distances between PACS domains and their nearest neighbors from Laion-400M, as shown in Figure 7. Our results indicate that target domains are, in fact, more similar to the pre-training data than source domains. We inspect nearest neighbors manually, and find even exact duplicates (Figure 6). Interestingly, while we found not only domain-level duplicates but also exact matches in the pre-training data, the Image Similarity Hypothesis is ultimately less predictive than the Alignment Hypothesis.

### v3

@@ 5 Experiments
We evaluate Domain Generalization (DG) performance across both pretraining-aligned (IP) and pretraining-misaligned (OOP) data. We adhere to the DomainBed evaluation methodology, where one domain is chosen as the target, and the remaining act as source domains. To maintain a sufficient amount of training data, we train all DG methods on the original, unsplit datasets. We use hyper-parameter values recommended by the original implementation authors for each method.
@@
After training, models are evaluated separately on IP and OOP subsets, as well as the original, unsplit test domain. This allows us to measure how well each method generalizes to both pretraining-aligned and pretraining-misaligned data. We follow the literature’s standard practice of computing performance per target domain and averaging the results across all domains. This method ensures that any domain imbalances do not disproportionately influence the final performance metrics.
5.1 Algorithms
Explicit Regularization towards Pretraining: Several recent DG methods leverage explicit regularization towards the initialization. These methods generally operate either in weight space (by regularizing or freezing model parameters) or in feature space (by aligning internal feature representations with those of the pretrained model). MIRO (Cha et al., 2022) minimizes the Mutual Information between DG model intermediate features and CLIP intermediate features. Attention Tuning (Teterwak et al., 2023; Touvron et al., 2022) freezes all parameters except those in the Multiheaded-Attention Layers. VL2V-SD (Addepalli et al., 2024) self-distills a linear combination of CLIP vision and text outputs into a model. CLIPood (Shu et al., 2023) regularizes both weights and output features. Weights are averaged between the pre-trained CLIP model and the fine-tuned model, and outputs are regularized using a loss that incorporates information from the pre-trained text encoder. Linear Probe - Fine Tuning (LP-FT) (Kumar et al., 2021) freezes the backbone, trains a linear probe, and then performs full finetuning. An untrained linear probe can cause finetuning to update the frozen backbone needlessly, potentially unlearning discriminative features. This biases the model towards pre-trained weights with smaller gradient updates.
Flat Optima: Studies by  Izmailov et al. (2018) and  Cha et al. (2021) have shown that flat minima generalize better than sharp minima, as they make loss values less sensitive to perturbations in the loss surface, resulting in smaller increases in loss during domain shifts. SWAD (Cha et al., 2021) Averages model parameters during training, determining the interval over which to average using validation loss over source domain. Model Parameter Averaging (MPA) (Arpit et al., 2022) Starts averaging model parameters after a number of burn-in steps to find flat minima. SAGM (Wang et al., 2023) is an optimizer that explicitly optimizes for flat minima.
Baseline and Oracle: We also train a baseline and oracle model for lower-bound and upper-bound reference. The baseline model is an Empirical Risk Minimization (ERM) model that is finetuned on source domains and evaluated on target domains, and has been found to be effective for the DG task (Gulrajani & Lopez-Paz, 2020). The oracle model is trained on an 80% training split of all domains and evaluated on a 20% test split. The oracle model removes the OOD aspect of generalization and provides a reasonable upper bound for DG methods. Finally, we add a zero-shot LAION-400M (Schuhmann et al., 2021) pre-trained ViT-B/16 from OpenClip, which is the pre-trained model used in our analysis.
5.2 Results
Table 1(c) reports results for the poorly pre-training aligned DomainBed-OOP, higly pre-training aligned DomainBed-IP, and standard DomainBed datasets. Underlined results highlight the best performance of any single method (excluding method combinations), while the bold numbers show the highest performance overall, excluding the upper bound and zeros-hot baseline. Overall, we find that AlignmentScore is highly predictive of performance. More detailed observations about the results follow, while a comparison to PerceptualSimilarityScore can be found in Appendix B.5.
DG methods perform well on DomainBed-IP: In most datasets within DomainBed-IP, domain generalization approaches achieve excellent performance. On two out of five datasets ( OfficeHome and PACS), the best DG method even outperforms the oracle. A notable exception is TerraIncognita, where CLIPood scores 30% below the oracle, highlighting this dataset as a challenging outlier. Interestingly, all three datasets where performance exceeds the oracle have an average IP AlignmentScore of 0.28, while the others have a lower average AlignmentScore of around 0.26. Therefore, the underperformance of TerraIncognita may be partially explained by its lower IP AlignmentScore, suggesting that alignment plays a significant role in DG performance, even within the IP case. Another interesting observation is that on the IP subset, the zero-shot model can achieve performances greater than the finetuned models (for PACS and VLCS). This means that, for IP-data, DG finetuning sometimes causes more catastrophic forgetting than learning of new features We show additional analysis of zero-shot models in Appendix B.9.
DG Methods leave much to be desired on DomainBed-OOP, but are still stronger than ERM: In DomainBed-OOP, we observe that even the top-performing DG methods struggle with low-alignment data. For example, CLIPood achieves 57.1% accuracy, which is a significant drop compared to its performance on DomainBed-IP (84.7%) and DomainBed-All (78.1%). Despite this, DG methods still outperform Empirical Risk Minimization (ERM) by up to 10% on DomainBed-OOP. Thus, DG methods are better equipped to handle domain shifts than ERM, possibly due to weak transfer of knowlegdge from pre-trained features. Nevertheless, there is still substantial room for improvement on low-alignment samples.
Table 1: Benchmarking DG methods on DB-IP/OOP. Domain generalization methods excel on high-alignment (IP) datasets—often even surpassing the oracle—while their performance noticeably drops on low-alignment (OOP) data, though still outperforming ERM.
@@ 5.2 Results
(a) Samples with high AlignmentScore values, indicating good pretraining alignment.
@@
SOTA methods do not consistently outperform older methods on OOP data: While CLIPood (Shu et al., 2023) clearly outperforms other methods on DomainBed-All, it performs comparably to older methods on DomainBed-OOP. For example, MIRO + MPA (Cha et al., 2022; Arpit et al., 2022) achieves 57.2% on DomainBed-OOP, which is nearly the same as CLIPood’s 57.1%. This suggests that CLIPood’s primary advantage comes from well-aligned samples in DomainBed-IP, where obtains 0.5% better performance than the next-best method.
@@
Model Parameter Averaging (MPA) boosts performance on OOP data: MPA obtains a significant 6% gain over ERM on DomainBed-OOP. MPA is 0.5% better than MIRO on OOP data despite MPA being 2% worse than MIRO on IP data. When combined with MIRO, MPA delivers the best performance on DomainBed-OOP, slightly surpassing CLIPood. This suggests that MPA can complement other regularization-based methods like MIRO. On DomainBed-IP, MIRO + MPA is less than 1% away from CLIPood’s performance, demonstrating versatility across both high- and low-alignment data. Interestingly, SWAD underperforms MPA on DomainBed-OOP by 2%, despite being conceptually similar. We attribute this to selecting the averaging interval on the source data, which introduces overfitting to source domains.
5.3 Discussion
As an increasing number of works in the Domain Generalization sub-field leverage pre-trained CLIP models for Domain Generalization benchmarks, it is important to better characterize the impacts of pre-training on DG. We leave the reader with the following takeaways:
Pre-training Alignment Predicts DG Performance: Our study demonstrates that pre-training alignment, measured as the cosine similarity between image and text embeddings, is a robust predictor of DG performance. This holds true even after source fine-tuning, highlighting that the quality of alignment achieved during pre-training has a significant impact on the generalization capability.
Current DG Methods Exploit Pre-training Rather Than Learning New Features: Our findings reveal a large difference in the performance of DG methods between pretraining-aligned (IP) and pretraining-misaligned (OOP) data. While state-of-the-art methods achieve near-oracle performance on IP data, they struggle significantly on OOP data. This indicates that current methods primarily leverage on pre-trained features rather than learning new, generalizable features from source data. Consequently, their success is heavily tied to the quality of pre-training, rather than fine-tuning.
Benchmarks Should Reflect Pre-training Reliance: The reliance on pre-trained alignment calls for a reevaluation of DG benchmarks. Existing benchmarks often aggregate results across all target data, masking the limitations of DG methods on low-alignment samples. To address this, we propose splitting evaluation datasets into In-Pretraining (IP) and Out-of-Pretraining (OOP) subsets. This provides a clearer picture of where DG methods succeed and where they fail. We hope that our proposed DomainBed-IP/OOP splits will guide the development of future methods that are better equipped to handle low-alignment data while maintaining performance on high-alignment samples.
@@ 6 Conclusion
We systematically explore how Domain Generalization (DG) methods rely on pre-trained feature alignment from models like CLIP. We hypothesize that the alignment between image and text embeddings during pre-training strongly predicts DG performance. Our experiments confirm this, showing that methods perform well on high-alignment samples (DomainBed-IP) but struggle on low-alignment data (DomainBed-OOP). Notably, state-of-the-art methods like CLIPood perform near oracle-level on aligned data but see significant drops on misaligned samples. This suggests current DG methods rely on pre-trained features and fail to learn new, generalizable features from source domains. Moving forward, two paths emerge: developing DG methods that better learn generalizable features, or focusing on improving pre-trained backbones. While foundation models will continue to advance, there will always be specialized distributions where they fail. Therefore we take the stance that improved DG finetuning remains an important avenure of research. We hope these findings inspire further research into improving generalization on low-alignment data, pushing DG beyond reliance on pre-trained alignment.
7 Acknowledgments
@@ A.1 Training and Evaluation Details
We use a slightly modified MIRO (Cha et al., 2022) codebase for training and evaluation. We use leave-one-out evaluation, where a model is trained on all domains except the evaluation domain. We emphasize that we use DomainBed-IP and DomainBed-OOP as evaluation data only, models are trained on full datasets.
For training, we use an OpenCLIP-ViT-B/16 (Ilharco et al., 2021) trained on LAION-400M (Schuhmann et al., 2021). We use default hyper-parameters as defined by  (Cha et al., 2022). This includes a learning rate of 5e-5, weight decay of 0.0, a batch size of 32 per-domain, an Adam Optimizer, and no dropout for all methods.
@@
For benchmarked methods, we also use hyper-parameters found to be best in respective papers. For SWAD, we use an optimum patience parameter value of 3, overfit patience parameter value of 6, and tolerance ratio of 6. For MIRO, we use use regularizer loss weight of 1.0. For CORAL, we use a CORAL regularizer weight of 1.0, following  (Cha et al., 2021). For LP-FT, we train the linear probe for 600 steps before unlocking the full backbone. For Model Parameter Averaging, we burn in the training for 600 steps before averaging iterates. For VL2V-SD and CLIPood, we directly use the author’s implementation and hyper-parameters, except initializing with OpenCLIP (Ilharco et al., 2021).
@@ B.1 AlignmentScore vs Accuracy
In Figure 11, we plot all benchmarked methods from the main paper, with x-axis corresponding to AlignmentScore, and the y-axis corresponding to the Top-1 Accuracy. We normalize for dataset size, so that no dataset dominates the count. In Figures 12 through 16, we plot these statistics independently per dataset, and find the trends consistent across datasets.
@@ B.2 Per-Dataset Benchmarking Results
We expand Table 1(c) in the main paper into per-dataset results in Table 10 through 24.
@@ B.3 Similarity of Target to Pre-training
To evaluate the Image Similarity Hypothesis, we retrieve the nearest neighbors from the Laion-400M dataset (Schuhmann et al., 2021). This raises the question of how similar the target domains are to the pre-training data and whether the source domains might be even more similar. To investigate this, we compute Maximum Mean Discrepancy (MMD) distances between PACS domains and their nearest neighbors from Laion-400M, as shown in Figure 5. Our results indicate that target domains are, in fact, more similar to the pre-training data than source domains. We inspect nearest neighbors manually, and find even exact duplicates (Figure 4). Interestingly, while we found not only domain-level duplicates but also exact matches in the pre-training data, the Image Similarity Hypothesis is ultimately less predictive than the Alignment Hypothesis.
@@ B.4 Other backbones
B.5 Splitting DomainBed using PerceptualSimilarityScore
In Figure 2 a.), we show that the slope of the relationship of Top-1 Accuracy vs Perceptual Similarity Score is positive but shallow. This suggests that using PerceptualSimilarityScore as an alternative to AlignmentScore for splitting DomainBed would not be very effective. To further prove this point, we split at a PerceptualSimilarityScore value of 0.86 in Table 3. We can see the differences are not very large between OOP and IP, indicating that AlignmentScore is a better thresholding metric.
Table 3: Splitting DomainBed by PerceptualSimilarityScore. The differences between IP and OOP with this split are much lower than with Alignment Score.
Dataset PACS VLCS TerraIncognita OfficeHome DomainNet
Perceptual IP 97.2 74.8 60.6 86.8 61.4
Perceptual OOP 95.6 77.4 42.9 78.3 55.3
B.6 Comparing AlignmentScore with zero-shot classification confidence score
We also consider an alignment score which takes into account uncertainties, and compute a score using the confidence of the zero-shot classifier formed by the pre-trained CLIP model for each sample. Specifically, for a sample with ground truth class c, we calculate the softmax over the logits output by the zero-shot classifier, and use the resulting probability p(c) as the score. We refer to this as the Calibrated AlignmentScore and show the results in Figure 6. Although the score predicts generalization for both OfficeHome and DomainNet, the scores have different scales for different datasets. In contrast out AlignmentScore does align across datasets to a greater degree (Figure 7)
Figure 6: Top-1 DG Accuracy vs calibrated alignment: We use the confidence of the zero-shot classifier formed by the pre-trained CLIP models an alignment measure. Although the score predicts generalization for both OfficeHome and DomainNet, the scores have different scales for different datasets.
Figure 7: Top-1 DG Accuracy vs AlignmentScore: The AlignmentScore introduced in our work have scales which are comparable across different datasets.
We explore the effect of averaging PerceptualSimilarityScore and AlignmentScore in Figure 8. We can see that there is not much of a compositional effect, so therefore we stick with AlignmentScore as our generalization predictor.
Figure 8: Combining PerceptualSimilarity Score and AlignmentScore: We explore the effect of averaging PerceptualSimilarityScore and AlignmentScore. There is no visible additional signal from averaging. We therefore stick with AlignmentScore as our generalization predictor.
B.7 Image Similarity Hypothesis for source data
The main drawback of the Image Similarity hypothesis is that it does not consider how well the nearest perceptual neighbor is learned during pre-training. One reason for a sample being poorly learned during pre-training is that the pre-training caption is not very relevant to the DG task. Source data is unlikely to have this issue, since source and target domains share labels. Therefore it is interesting to ask how strongly correlated the PerceptualSimilarityScore is with DG accuracy when measured between source and target. Indeed , as seen in Figure 9, there is a strong correlation. However, simply using source-data to compute the PerceptualSimilarity results in an incomplete understanding of the relationship between target data and the training procedure, due to the lack of consideration of the pre-training. In fact, zero-shot models with NO learning from source are very performant (Table 1(c))
Figure 9: PerceptualSimilarity to Source for OfficeHome. DG performance is correlated with similarity to source images.
B.8 Training from scratch
We focus on analyzing the generalization capabilities of CLIP-based DG methods, as these have been demonstrated to be the strongest in recent work. However,from-scratch experiments are valuable to measure how much finetuning on source data can learn, and so we also trained ResNet-50 models from scratch, using the MPA (Model Parameter Averaging) method and present the results in Table 4. We can see that performance is very low even on ALL DomainBed data (unsplit), further confirming that existing DG algorithms are poor at learning new knowledge from source data alone, and they rely on strong pre-training for good performance.
Table 4: Training ResNet-50 from scratch. DG performance is very poor, showing DG methods rely on strong pre-training
DomainNet OfficeHome PACS TI VLCS Avg
ResNet-50 from Scratch 7.6 17.2 36.3 23.8 55.2 28.0
Clip ViT-B/16 58.9 82.0 94.3 50.7 82.3 73.6
B.9 Analysis of Zero-Shot Behavior
To better understand the behavior of Domain Generalization (DG) methods, we compute confusion matrices for zero-shot (ZS) and In-Pretraining (IP) and Out-of-Pretraining (OOP) subsets for the PACS dataset using CLIPood. The results are shown in Tables 5 and 6.
Table 5: Confusion matrix for the IP subset of the PACS dataset.
DG: Incorrect DG: Correct
ZS: Incorrect 141 70
ZS: Correct 179 8231
Table 6: Confusion matrix for the OOP subset of the PACS dataset.
DG: Incorrect DG: Correct
ZS: Incorrect 98 62
ZS: Correct 15 890
The confusion matrices reveal key insights about the behavior of DG methods:
  • For the IP subset, DG methods catastrophically forget a significant number of samples, flipping correct predictions made by zero-shot models to incorrect ones. This suggests that DG methods can even have a negative value for IP data. Nevertheless, performance remains strong.
  • For the OOP subset, DG methods flip very few correct samples to incorrect ones. However, the state-of-the-art (SOTA) method, CLIPood, is still unable to correct the majority of incorrectly classified samples in this subset.
Overall, in both IP and OOP cases, the results indicate that CLIPood and other DG methods rely heavily on pre-training. These findings underscore the limitations of current DG methods and highlight areas for improvement.